[Brahms-l] Shift report 20040213 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Feb 13 2004 - 07:44:27 EST
------------------------------------------------------------
Date: Fri Feb 13 07:44:27 2004, Shift: 00:00-08:00
Supervisor: Ionut ARSENE, with: Radek KARABOVICZ
Log book#15 pages 55 - 55

Shift Summary:
------------------------------------------------------------
1. We had beam for 6 hours.

2. Data runs achieved: 10309 - 10313
 We finished the B861 in FS at 4 degrees setting and started 
the B2442 setting. MRS is still at 45 degrees with D5 at 
B2255.
 We have taken ~850k trig2 events and ~750k trig3 events.
 10307 and 10308 are pedestal runs.
 
3. The D2 current at 1/4 field is not stable. Is oscillating 
with +/-30 Amps around the nominal value.
  One mouse is dead... behind the fridge. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 13 07:44:45 2004

This archive was generated by hypermail 2.1.8 : Fri Feb 13 2004 - 07:45:05 EST